# On the Generality and Persistence of Cosmological Stasis

### v1

@@ I Introduction
One of the hallmark characteristics of an expanding universe is the time-evolution of components that contribute to the overall energy density of the universe. It has recently been introduced [1] that this assumption is not always true, and in fact the universe could experience extended phases of cosmological “stasis” in which the cosmological abundances of matter, radiation, and/or vacuum energy can remain steady over extended $e$-folds of cosmological evolution, facilitated by various physical mechanisms driving energy pumps which oppose the effects of cosmological expansion [2]. These phenomena can arise naturally from a variety of beyond the Standard Model (BSM) physical theories, for instance in those that predict towers of unstable states that inevitably decay into other possibly mixed-component states.
@@ II Maximizing Stasis with Differentiable Simulations
For $N$ species, our task is to solve a set of $N+1$ coupled ODEs as given in equations 4 and 8 and compute the number of $e$-folds of a potential stasis epoch. We utilize diffrax [9], a jax-based [10] library that provides numerical differential equation solvers that preserve gradients and also allow backpropagation through our solutions to the Boltzmann equations. jax is a differentiable high-performance numerical computing library which utilizes just-in-time (JIT) compilation and vectorized computations, earning popularity in the sciences for its efficiency and modularity. The set of ODEs is stiff, meaning certain numerical techniques will be unstable without a sufficiently small resolution for the step-size. We use the Kvaerno5 [11] solver which is a 5th order explicit singly diagonal implicit Runge-Kutta (ESDIRK) method suited for stiff ODEs, with an absolute and relative tolerance $\text{atol}=\text{rtol}=10^{-8}$ in the step size. We terminate the solver at $t=t_{max}$ when $\Omega_{M}(t)=10^{-4}$, indicating that the universe has passed into a radiation-dominated epoch. Backpropagation through a differential equation can also be very memory expensive. We use RecursiveCheckpointAdjoint [12] which utilizes a binomial checkpointing scheme to preserve memory usage during backpropagation.
@@ II.1 $\epsilon$-Stasis
This differentiable calculation, however, requires that $\overline{\Omega}_{M}$ is known. This is retrieved from a separate, differentiable abundance-finder algorithm. The algorithm first marks $\Omega_{M}$ values that are not matter dominated or radiation dominated. It then determines links between consecutive valid values that are within a specified tolerance $\epsilon$. It proceeds to count these links for each time step and identifies the range of indices with the maximum number of links, indicating a period of stasis. Finally, it filters the valid $\Omega_{\text{M}}$ values within this range and computes $\overline{\Omega}_{M}$ as the median of these values, representing the typical abundance during the stasis period. This algorithm, of course, also yields a prediction for $\mathcal{N}$, but we find by examining solutions that it is inaccurate and does not function as well as equation II.1.
@@ II.2 Maximizing Stasis
Sort-correlation may be performed with any off-the-shelf sorting algorithm, including but not limited to a simple jax.numpy.sort. As $\mathrm{\Gamma_{\ell}}$ and $\mathrm{\Omega_{\ell}^{(0)}}$ are updated with gradient ascent, we would like to continue to enforce that their spectra are sort-correlated. Doing so crucially requires that they are sorted differentiably, to not interrupt the flow of gradients during optimization. For this reason, we implement a differentiable bitonic sorting algorithm inspired by [13]. This algorithm works by recursively dividing an array into smaller sub-arrays, sorting them, and then merging them using compare-and-swap operations, the latter of which is modified to be differentiable. This sorting technique is also used when doing SVI experiments.
@@
for a given step $i$, where $\eta(t)$ is a (potentially) time-dependent learning rate. Optimization is subject to an early-stopping criterion if $f(\theta)$ does not improve over a specified number of epochs $\xi$ or if a NaN is encountered during optimization.
@@
Figure 3: Experiments showing a preference for the exponential model upon optimizing stasis with gradients for uniform and power-law initializations for $N=50$ and $\Gamma_{N-1}/H^{(0)}=0.1$. Gradient ascent was subject to the constraint $0.80.2<\overline{\Omega}_{M}<$ 0.8. Optimization was done for 50 random initializations for 50,000 epochs with early-stopping. An initial learning rate of .01 was used with 0.1 multiplicative decay at epoch 10,000. A clear bias towards an exponential model is shown for optimal $\Gamma_{\ell}$ and $\Omega_{\ell}^{(0)}$, even when initialized with a power-law distribution similar to the original model of stasis. It is also seen that the drift shifts across the 1:1 line dividing exponential and power-law confidence equality. In some instances, the initialized parameters are shown to already be a good exponential fit due to the compression of the relative abundance and decay spectra. Even under such conditions, the drift towards a more exponential model and away from a power-law model is evident.
@@ II.3 Stasis and Unsorted Abundances
The result of this experiment is shown in Figure 4. Upon convergence, we find that the gradient ascent algorithm has learned to only sort the species contributing to stasis; that is, robust stasis epochs are possible without strictly correlated decay rates and abundances. This realization would have been difficult to make with analytical modeling!
@@ IV.1 The Evidence Lower Bound
for two continuous distributions $P$ and $Q$. The KL divergence is a type of statistical “distance”, measuring how different the two distributions are. It is always positive-definite, satisfying $D_{KL}(P||Q)\geq 0$; however, it is not a metric in a formal sense as it is not a symmetric quantity (i.e. $D_{KL}(P||Q)\neq D_{KL}(Q||P))$.
@@ V.1 An Exact Exponential Model of Stasis
Motivated by our numerical results, we introduce an exponential model of stasis that yields parametrically more stasis than power-law models, as a function of the number of species. We will show that such a model obeys the dynamical equations of stasis and will result in robust, finite epochs of stasis. We additionally derive algebraic relations with exact predictions for $\mathcal{N}$ and $\overline{\Omega}_{M}$ from our model.
@@
With this result in hand, we can proceed to compare coefficients in 15 and 59, in which we find
@@
We see that the exponential model of stasis yields parametrically more e-folds than the power-law model, which exhibited $\mathcal{N}\sim\log(N)$ scaling. See Figure 10 for a comparison of stasis plots with $N=300$ species in the exponential and power-law model. A numerical comparison of $e$-fold data and the theoretical prediction for $\gamma=1$, $\alpha=2/7$, and $\Gamma_{N}=0.01$ yielding a stasis epoch with matter-radiation equality (MRE), is shown in Figure 9. As expected, there is excellent agreement between data and theory in the limit $N\rightarrow\infty$. The figure also illustrates that more than 60 $e$-folds of MRE can be achieved with just 100 species in the exponential model. This result shows that a relatively small number of species is needed to achieve inflation-level $e$-folding, which may have important phenomenological implications.
@@ VI Discussion
A stasis epoch could have significant cosmological implications [1], potentially affecting dark matter production, large-scale structure formation, and the overall understanding of cosmological evolution and the age of the Universe. Of particular interest for $M\rightarrow\gamma$ stasis is the possibility of stasis epochs occurring after a matter-dominated epoch at the end of inflation or after a later matter-dominated epoch prior to nucleosynthesis. In the first case, the decay of the inflaton can instead source $\phi_{\ell}$ states. The hierarchical decays of those states during stasis would then be the source of inflationary reheating, with the universe entering RD with the conclusion of the decays. In the second case, the tower of matter fields can lead to an early matter dominated epoch prior to BBN, but well after inflation, the decays of which give rise to stasis. There are a number of potential phenomenological implications of such an epoch, but a stasis state is an attractor regardless of the arena in which $\phi_{\ell}$ are produced.
@@
Our methodology began with constructing a differentiable Boltzmann solver, which is the crux of our approach. We maximized the number of $e$-folds of stasis by following gradients in our differentiable simulation with model-agnostic (aside from the prior) random initializations of $\mathrm{\Gamma_{\ell}}$ and $\mathrm{\Omega_{\ell}^{(0)}}$. Such stasis-maximizing trajectories through the space of rates and abundances motivated the study of log-uniform distributions, which we compared to power-law distributions by taking random samples and solving the Boltzmann equations. This comparison provided an initial indication of the significant discrepancy in the duration of stasis epochs between the two models. With distributional priors established, we employed SVI with normalizing flows for a Bayesian analysis of the full, high-dimensional parameter space; there are some essential differences from standard SVI that we discuss in the text. From the posteriors, we again saw that stasis maximization generally prefers log-uniform distributed $\mathrm{\Omega_{\ell}^{(0)}}$ and $\mathrm{\Gamma_{\ell}}$, corresponding to a model that is exponential in the species index $\ell$. It was also observed that a qualitative hybrid of power-law and exponential parameters in the posterior could result in epochs of stasis. Since this motivates a mixed model, it further indicates the generality is stasis.
@@
  • Stasis epochs can arise without strictly correlated rates and abundances, especially when some species do not participate in the stasis dynamics.

### v3

@@ I Introduction
One of the hallmark characteristics of an expanding universe is the time-evolution of components that contribute to the overall energy density of the universe. It has recently been introduced [1] that this assumption is not always true, and in fact the universe could experience extended phases of cosmological “stasis” in which the cosmological abundances of matter, radiation, and/or vacuum energy can remain steady over extended $e$-folds of cosmological evolution, facilitated by various physical mechanisms driving energy pumps which oppose the effects of cosmological expansion [2]. These phenomena can arise naturally from a variety of beyond the Standard Model (BSM) physical theories, for instance in those that predict towers of unstable states that inevitably decay.
@@ II Maximizing Stasis with Differentiable Simulations
For $N$ species, our task is to solve a set of $N+1$ coupled ODEs as given in equations 4 and 8 and compute the number of $e$-folds of a potential stasis epoch. We utilize diffrax [9], a jax-based [10] library that provides numerical differential equation solvers that preserve gradients and also allow backpropagation through our solutions to the Boltzmann equations. jax is a differentiable high-performance numerical computing library which utilizes just-in-time (JIT) compilation and vectorized computations, earning popularity in the sciences for its efficiency and modularity. The set of ODEs is stiff, meaning certain numerical techniques will be unstable without a sufficiently small resolution for the step-size. We use the Kvaerno5 [11] solver which is a 5th order explicit singly diagonal implicit Runge-Kutta (ESDIRK) method suited for stiff ODEs, with an absolute and relative tolerance $\text{atol}=\text{rtol}=10^{-8}$ in the step size. We terminate the solver at $t=t_{max}$ when $\Omega_{M}(t)=10^{-4}$, indicating that the universe has passed into a radiation-dominated epoch. Backpropagation through a differential equation can also be very memory expensive. We use RecursiveCheckpointAdjoint [12, 13] which utilizes a binomial checkpointing scheme to preserve memory usage during backpropagation.
@@ II.1 $\epsilon$-Stasis
This differentiable calculation, however, requires that $\overline{\Omega}_{M}$ is known. This is retrieved from a separate, differentiable abundance-finder algorithm. The algorithm first marks $\Omega_{M}$ values that are not matter dominated or radiation dominated (i.e. $0.990.01<\Omega_{M}<$ 0.99). It then determines links between consecutive valid values that are within a specified tolerance $\epsilon$. It proceeds to count these links for each time step and identifies the range of indices with the maximum number of links, indicating a period of stasis. Finally, it filters the valid $\Omega_{\text{M}}$ values within this range and computes $\overline{\Omega}_{M}$ as the median of these values, representing the typical abundance during the stasis period. This algorithm, of course, also yields a prediction for $\mathcal{N}$, but we find by examining solutions that it is inaccurate and does not function as well as equation II.1.
@@ II.2 Maximizing Stasis
Sort-correlation may be performed with any off-the-shelf sorting algorithm, including but not limited to a simple jax.numpy.sort. As $\mathrm{\Gamma_{\ell}}$ and $\mathrm{\Omega_{\ell}^{(0)}}$ are updated with gradient ascent, we would like to continue to enforce that their spectra are sort-correlated. Doing so crucially requires that they are sorted smoothly and differentiably, to not interrupt the flow of gradients during optimization. For this reason, we implement a custom differentiable bitonic sorting algorithm inspired by [14]. This algorithm works by recursively dividing an array into smaller sub-arrays, sorting them, and then merging them using compare-and-swap operations, the latter of which is modified to be differentiable. This sorting technique is also used when doing SVI experiments.
@@
for a given step $i$, where $\eta(t)$ is a (potentially) time-dependent learning rate. In subsequent experiments, a decay-factor $\gamma$ is applied at epoch $t^{\prime}$ such that the learning rate has the functional form
$\eta(t)=\begin{cases}\eta_{0},&\text{for }t<t^{\prime}\\ \gamma\cdot\eta_{0},&\text{for }t\geq t^{\prime}\end{cases}$ (26)
Optimization is subject to an early-stopping criterion if $f(\theta)$ does not improve over a specified number of epochs $\xi$ or if a NaN is encountered during optimization.
@@
Figure 3: Experiments showing a preference for the exponential model upon optimizing stasis with gradients for uniform and power-law initializations for $N=50$ and $\Gamma_{N-1}/H^{(0)}=0.1$. Gradient ascent was subject to the constraint $0.80.2<\overline{\Omega}_{M}<$ 0.8. Optimization was done for 50 random initializations for 50,000 epochs with early-stopping. An initial learning rate $\eta_{0}=0.01$ was used with a $\gamma=0.1$ multiplicative decay at epoch $t^{\prime}=10,000$. A clear bias towards an exponential model is shown for optimal $\Gamma_{\ell}$ and $\Omega_{\ell}^{(0)}$, even when initialized with a power-law distribution similar to the original model of stasis. It is also seen that the drift shifts across the 1:1 line dividing exponential and power-law confidence equality. In some instances, the initialized parameters are shown to already be a good exponential fit due to the compression of the relative abundance and decay spectra. Even under such conditions, the drift towards a more exponential model and away from a power-law model is evident.
@@ II.3 Stasis and Unsorted Abundances
The result of this experiment is shown in Figure 4. This analysis has indeed shown that robust stasis epochs are possible without strictly correlated decay rates and abundances, even though the species contributing to stasis have abundances increasing with decay rates. This realization would have been difficult to make with analytical modeling!
@@ IV.1 The Evidence Lower Bound
for two continuous distributions $P$ and $Q$. The KL divergence is a type of statistical “distance”, measuring how different the two distributions are. It is always positive semi-definite, satisfying $D_{KL}(P||Q)\geq 0$; however, it is not a metric in a formal sense as it is not a symmetric quantity (i.e. $D_{KL}(P||Q)\neq D_{KL}(Q||P))$.
@@ V.1 An Exact Exponential Model of Stasis
Motivated by our numerical results, we introduce an exponential model of stasis that yields parametrically more stasis than power-law models, as a function of the number of species. We will show that such a model obeys the dynamical equations of stasis and will result in robust, finite epochs of stasis. We additionally derive algebraic relations with exact predictions for $\mathcal{N}$ and $\overline{\Omega}_{M}$ from our model.i
@@
With this result in hand, we can proceed to compare coefficients in 15 and 60, in which we find
@@
We see that the exponential model of stasis yields parametrically more e-folds than the power-law model, which exhibited $\mathcal{N}\sim\log(N)$ scaling. A numerical comparison of $e$-fold data and the theoretical prediction for $\gamma=1$, $\alpha=2/7$, and $\Gamma_{N}=0.01$ yielding a stasis epoch with matter-radiation equality (MRE), is shown in Figure 9. Further, see Figure 10 for a comparison of stasis plots with $N=300$ species in the exponential and power-law model. As expected, there is excellent agreement between data and theory in the limit $N\rightarrow\infty$. The figure also illustrates that more than 60 $e$-folds of MRE can be achieved with just 100 species in the exponential model. This result shows that a relatively small number of species is needed to achieve inflation-level $e$-folding, which may have important phenomenological implications.
@@ VI Discussion
A stasis epoch could have significant cosmological implications [1], potentially affecting dark matter production, large-scale structure formation, and the overall understanding of cosmological evolution and the age of the Universe. Of particular interest for $M\rightarrow\gamma$ stasis is the possibility of stasis epochs occurring after a matter-dominated epoch at the end of inflation or after a later matter-dominated epoch prior to nucleosynthesis. In the first case, the decay of the inflaton can instead source $\phi_{\ell}$ states. The hierarchical decays of those states during stasis would then be the source of inflationary reheating, with the universe entering radiation domination with the conclusion of the decays. In the second case, the tower of matter fields can lead to an early matter dominated epoch prior to BBN, but well after inflation, the decays of which give rise to stasis. There are a number of potential phenomenological implications of such an epoch, but a stasis state is an attractor regardless of the arena in which $\phi_{\ell}$ are produced.
@@
Our methodology began with constructing a differentiable Boltzmann solver, which is the crux of our approach. We maximized the number of $e$-folds of stasis by following gradients in our differentiable simulation with model-agnostic (aside from the prior) random initializations of $\mathrm{\Gamma_{\ell}}$ and $\mathrm{\Omega_{\ell}^{(0)}}$. Such stasis-maximizing trajectories through the space of rates and abundances motivated the study of log-uniform distributions, which we compared to power-law distributions by taking random samples and solving the Boltzmann equations. This comparison provided an initial indication of the significant discrepancy in the duration of stasis epochs between the two models. With distributional priors established, we employed SVI with normalizing flows for a Bayesian analysis of the full, high-dimensional parameter space; there are some essential differences from standard SVI that we discuss in the text. From the posteriors, we again saw that stasis maximization generally prefers log-uniform distributed $\mathrm{\Omega_{\ell}^{(0)}}$ and $\mathrm{\Gamma_{\ell}}$, corresponding to a model that is exponential in the species index $\ell$. This motivates a hybrid power-law / exponential model beyond the strict power-law or exponential models we have studied, further indicating the generality of stasis.
@@
  • Stasis epochs can arise without strictly correlated rates and abundances, with the species participating in stasis having abundances increasing with decay rates.